# KERMIT: Knowledge Graph Completion of Enhanced Relation Modeling with Inverse Transformation

### v2

@@ 4.2. Model architecture
(14) $X_{hr}=[x_{\text{CLS}},H_{\text{desc}},x_{\text{SEP}},R_{\text{desc}},x_{\text{SEP}},T_{\text{pred}},x_{\text{SEP}}]$
Following SimKGC (Wang et al., 2022), we utilize two BERT (Devlin et al., 2019) encoders with identical pre-trained parameters, operating independently during training and inference for different representation learning purposes.
Given a triple $(h,r,t)$, the encoded vectors of $h,r$ and $t$ are obtained via the BERT encoders following a pooling strategy:
(15) $e_{hr}$ $=\text{Pool}(\text{BERT}_{hr}(X_{hr}))$
(16) $e_{t}$ $=\text{Pool}(\text{BERT}_{t}(X_{t}))$
Regarding the pooling method, we deviate from the prevalent use of the CLS token. Instead, we opt for mean pooling, as it has been demonstrated to yield superior sentence embeddings (Reimers and Gurevych, 2019).
4.3. Supervised contrastive training
In order to enhance the expressiveness of query and entity representations, we propose to effectively leverage the label information inherent in a KG, thus formulating a supervised variant (Khosla et al., 2020) of the self-supervised contrastive loss shown in Equation 8:
(17) $\mathcal{L}=-\frac{1}{|\mathcal{N}^{+}|}\sum_{t\in\mathcal{N}^{+}}\log\frac{e^{(f(h,r,t)-\gamma)/\tau}}{e^{(f(h,r,t)-\gamma)/\tau}+\sum_{t^{\prime}\in\mathcal{N}^{-}}{e^{(f(h,r,t^{\prime})-\gamma)/\tau}}}$
@@ 4.3. Supervised contrastive training
For sampling positive and negative entities, we utilize an in-batch strategy (Chen et al., 2020; Khosla et al., 2020) for the sake of efficiency. In this approach, class labels are assigned to entities within the same batch to indicate their positivity or negativity. Exploring more effective sampling strategies is reserved for future work. For instance, an entity’s embedding can be calculated using its textual description through a pre-trained language model, and entities with the lowest semantic similarity can be considered negatives.
Note that the predictive descriptions (i.e. $t_{\text{pred}}$ and $h_{\text{pred}}$) and inverse relations for all training and testing samples are obtained within a number of chat completion API invocations in advance, which is not part of the training and inference flow. An overview of our model is shown in Figure 2.
5. Experiments
5.1. Experiment setting
Table 3. Statistics of datasets.
Dataset Relation Entity Train Validation Test
@@ 5.1. Experiment setting
Table 4. Knowledge graph completion results for and WN18RR and FB15k-237 datasets. All metrics are in %. The best performance for each metric within each dataset is highlighted in bold, and the best metrics among categories are underlined.
WN18RR FB15k-237
MRR Hit@1 Hit@3 Hit@10 MRR Hit@1 Hit@3 Hit@10
NeuralLP (Yang et al., 2017) 38.1 36.8 38.6 40.8 23.7 17.3 25.9 36.1
DRUM (Sadeghian et al., 2019) 38.2 36.9 38.8 41.0 23.8 17.4 26.1 36.4
RNNLogic (Qu et al., 2020) 51.3 47.1 53.2 59.7 34.9 25.8 38.5 53.3
LERP (Han et al., 2023) 62.2 59.3 63.4 68.2 - - - -
TransE (Bordes et al., 2013) 24.3 4.3 44.1 53.2 27.9 19.8 37.6 44.1
DistMult (Yang et al., 2014) 44.4 41.2 47.0 50.4 28.1 19.9 30.1 44.6
RotatE (Sun et al., 2019) 47.6 42.8 49.2 57.1 33.8 24.1 37.5 53.3
ConvE (Dettmers et al., 2018) 43.0 40.0 44.0 52.0 32.5 23.7 35.6 50.1
TuckER (Balazevic et al., 2019) 47.0 44.3 48.2 52.6 35.8 26.6 39.4 54.4
KG-BERT (Yao et al., 2019) 21.6 4.1 30.2 52.4 - - - 42.0
MTL-KGC (Kim et al., 2020) 33.1 20.3 38.3 59.7 26.7 17.2 29.8 45.8
StAR (Wang et al., 2021b) 40.1 24.3 49.1 70.9 29.6 20.5 32.2 48.2
SimKGC (Wang et al., 2022) 66.6 58.7 71.7 80.0 33.6 24.9 36.2 51.1
CP-KGC (Zhang et al., 2023) 67.3 59.9 72.1 80.4 33.8 25.1 36.5 51.6
GHN (Qiao et al., 2023) 67.8 59.6 71.9 82.1 33.9 25.1 36.4 51.8
KERMIT (qwen2-7b) 69.4 62.0 73.7 83.3 35.1 25.8 38.7 53.7
KERMIT (phi3-small) 69.5 62.1 73.7 83.0 35.2 25.8 38.8 53.9
KERMIT (phi3-medium) 70.0 62.9 73.8 83.2 35.9 26.6 39.6 54.7
@@
Data preparation. To evaluate our methods for completing missing triples in KGs, we select two datasets for KGC task: WN18RR and FB15k-237 (Wang et al., 2021a). Statistics about each dataset are presented in Table 3. The WN18 and FB15k datasets were initially introduced in the work by (Bordes et al., 2013). Subsequent research (Toutanova et al., 2015; Dettmers et al., 2018), however, revealed that these two datasets are susceptible to test leakage. This issue led to the creation of refined versions: WN18RR and FB15k-237. The textual descriptions for WN18RR and FB15k-237 are sourced from the data made available by KG-BERT (Yao et al., 2019) and CP-KGC (Yang et al., 2024b), respectively.
@@
During the training process for the task of KGC, each triple $(h,r,t)$ is augmented by creating two queries: $(h,r,?)$ and $(t,r^{\prime},?)$, with corresponding answers $t$ and $h$. Here, $r^{\prime}$ represents the inverse of $r$. Since obtaining $h_{\text{pred}}$ in Equation 12 necessitates inverse relations, we utilize gpt4-turbo-2024-04-09 to generate these inverses, as outlined in Equation 13. The pairs of $r$ and $r^{\prime}$ for the WN18RR dataset are listed in Table 8.
@@
However, dealing with the FB15k-237 dataset appears more complicated. In contrast to the WN18RR dataset, the relations found within FB15k-237 pose a challenge due to their intractability for direct input into a text-based model. To address this concern, we undertake sentence-formation of the relations and identify their corresponding inverses with the aid of an LLM. Given the substantial number (237) of relations within the FB15k-237 dataset, we opt to present only a limited subset of these relations in Table 9.
@@
Evaluation Metrics. For each query $(h,r,?)$ or $(t,r^{\prime},?)$, our approach computes a score for each entity, along with determining the rank of the correct answer. The Mean Reciprocal Rank (MRR) and Hit@$k$ values under the filtered protocol (Bordes et al., 2013) are reported based on the computed ranks across all queries. MRR calculates the mean of the reciprocal ranks for the answer entities, while Hit@$k$ determines the percentage of desired entities ranked within the top $k$ positions. The filtered protocol excludes the scores of all known true triples in the training, validation, and testing sets.
@@
Implementation details. To obtain a comprehensive evaluation on the selection of LLMs, we utilize various open-source LLMs to generate predictive descriptions, including Qwen2-7B-Instruct (Yang et al., 2024a), phi-3-small-128k-instruct (Abdin et al., 2024) and phi-3-medium-4k-instruct (Abdin et al., 2024). In this paper, we refer to these three LLMs as qwen2-7b, phi3-small and phi3-medium for simplicity. A subset of triple examples of WN18RR with textual descriptions and predictive descriptions via phi3-medium is shown in Table 10 (more examples are shown in Table 12 and 11). For additional information, please see Appendix A.
@@
Table 5. Knowledge graph completion accuracy of KERMIT and its variants on WN18RR and FB15k-237. "pred-desc": with predictive descriptions; "few-shot": with few-shot triple examples when generating predictive descriptions; "sup": with supervised contrastive loss.
@@
Figure 3. Model performance of text-based methods with the proposed approaches on WN18RR dataset. "w/ inv. rel.": with only inverse relations; "w/ pred. desc": with predictive descriptions. The predictive descriptions are collected using phi3-medium.
5.2. Main results
In this section, we report empirical results of our method, along with those of the chosen algorithms. We adopt the numerical results for NeuralLP, DRUM, RNNLogic, and ConvE as presented in (Qu et al., 2020). For LERP, we refer to the values provided in (Han et al., 2023). The metrics for TransE, DistMult, RotatE, TuckER, KG-BERT, MTL-KGC, StAR, and SimKGC are sourced from (Wang et al., 2022; Zhang et al., 2023).
In Table 4, our proposed model exhibits state-of-the-art performance on the WN18RR and FB15k-237 datasets across all evaluation metrics. Compared to our baseline SimKGC, our model shows substantial improvements in MRR (rising from 66.6% to 70.0%), Hit@1 (increasing from 58.7% to 62.9%), and other metrics. Similar enhancements are observed on the FB15k-237 dataset.
5.3. Ablation results
@@
To further demonstrate the effectiveness and generalization ability of predictive descriptions and curated inverse relations, we apply our method to various text-based models, including SimKGC (Wang et al., 2022), StAR (Wang et al., 2021b), and KG-BERT (Yao et al., 2019), evaluating their performance on the WN18RR and FB15k-237 datasets (see Figure 3 and 4). The results show a distinct improvement across all metrics after incorporating predictive descriptions and inverse relations, demonstrating our method’s effectiveness as a general enhancement tool for text-based models.
Table 6. Results given by different tokenization strategies of the query sequence on the WN18RR and FB15k-237 datasets. "sep", "space" and "only $T_{\text{pred}}$" refer to the tokenization strategies specified by Equation 14, 18 and 19, respectively.
@@
Table 7. Predictive descriptions of the triple (cover, derivationally related form, covering) generated by phi3-medium, both with and without few-shot examples in the prompt. Descriptions of the entities involved are available in Table 10.
@@
6. Analysis
6.1. Fine-grained analysis
@@
(18) $X_{hr}=[x_{\text{CLS}},H_{\text{desc}},x_{\text{SEP}},R_{\text{desc}},x_{\text{SPACE}},T_{\text{pred}},x_{\text{SEP}}]$
However, since the operation complexity of multi-head attention (Vaswani et al., 2017) grows quadratically as the length of the input sequence increases (Shazeer, 2019), extra information might cause unexpected computational overhead. An alternative version is to represent $X_{hr}$ with just the $T_{\text{pred}}$ to reduce the computational cost, resulting in Equation 19.
(19) $X^{\prime}_{hr}=\left[x_{\text{CLS}},T_{\text{pred}},x_{\text{SEP}}\right]$
Few-shot triple examples play an important role in the prompting template we use to instruct LLMs, as illustrated in Table 1. We examine the impact of these few-shot examples by comparing the predictive descriptions generated by phi3-medium with and without their inclusion in the prompt, as shown in Table 7. When presented with the query (cover, derivationally related form, ?), incorporating few-shot examples clearly enhances the LLM’s understanding of the task. This results in a more precise and focused predictive description over the unknown answer ’covering’.
6.2. Case study on selected LLMs
@@ Appendix A Experiment settings
Comparison of algorithms. In experiment, the performance of our model is compared with that of the following algorithms: (1) Rule learning methods. We have selected a set of algorithms that specialize in learning logical rules for the purpose of completing missing triples. This selection allows for a comprehensive comparison, and the chosen algorithms are NeuralLP (Yang et al., 2017), DRUM (Sadeghian et al., 2019), RNNLogic (Qu et al., 2020) and LERP (Han et al., 2023). (2) Embedding-based methods. Furthermore, we extend our comparison to include various embedding methods, such as TransE (Bordes et al., 2013), DistMult (Yang et al., 2014), RotatE (Sun et al., 2019), ConvE (Dettmers et al., 2018) and TuckER (Balazevic et al., 2019). These models adopt an approach where representations for entities and relations are learned respectively. (3) Text-based methods. We select the following models that also leverage textual information into our consideration: KG-BERT (Yao et al., 2019), MTL-KGC (Kim et al., 2020), StAR (Wang et al., 2021b), SimKGC (Wang et al., 2022), CP-KGC (Zhang et al., 2023), GHN (Qiao et al., 2023).
Implementation details. To ensure a fair comparison with previous studies, we initialize the BERT encoders using pretrained weights from bert-base-uncased (English). The AdamW optimizer (Loshchilov and Hutter, 2019) with linear learning rate decay is employed to optimize the model parameters. The initial learning rates are set to $8\times 10^{-5}$ for the WN18RR dataset and $1\times 10^{-5}$ for the FB15k-237 dataset. We maintain a fixed batch size of 1024. The temperature parameter $\tau$ is tuned by searching within the set $\{0.02,0.03,0.04\}$, based on results from 65 epochs for WN18RR and 10 epochs for FB15k-237.

### v3

@@ 4.2 Model architecture
$X_{hr}=[x_{\text{CLS}},H_{\text{desc}},x_{\text{SEP}},R_{\text{desc}},x_{\text{SEP}},T_{\text{pred}},x_{\text{SEP}}]$ (12)
Following SimKGC [11], we utilize two BERT [40] encoders with identical pre-trained parameters, operating independently during training and inference for different representation learning purposes.
Given a triple $(h,r,t)$, the encoded vectors of $(h,r)$, and $t$, denoted as $e_{hr}$ and $e_{t}$, are computed following Equation 1 and 4. For the pooling method, we diverge from the commonly used CLS token and instead employ mean pooling, which has been shown to produce superior sentence embeddings [46]. An overview of our model is shown in Figure 2.
4.3 Supervised contrastive training
In order to enhance the expressiveness of query and entity representations, we propose to effectively leverage the label information inherent in a KG, thus formulating a supervised variant [13] of the self-supervised contrastive loss shown in Equation 6:
$\mathcal{L}=-\frac{1}{|\mathcal{N}^{+}|}\sum_{t\in\mathcal{N}^{+}}\log\frac{e^{(f(h,r,t)-\gamma)/\tau}}{e^{(f(h,r,t)-\gamma)/\tau}+\sum_{t^{\prime}\in\mathcal{N}^{-}}{e^{(f(h,r,t^{\prime})-\gamma)/\tau}}}$ (13)
@@ 4.3 Supervised contrastive training
For sampling positive and negative entities, we utilize an in-batch strategy [13, 25] for the sake of efficiency. In this approach, class labels are assigned to entities within the same batch to indicate their positivity or negativity. Exploring more effective sampling strategies is reserved for future work. For instance, an entity’s embedding can be calculated using its textual description through a pre-trained language model, and entities with the lowest semantic similarity can be considered negatives.
Note that the predictive descriptions (i.e. $t_{\text{pred}}$ and $h_{\text{pred}}$) and inverse relations for all training and testing samples are obtained within a number of chat completion API invocations in advance, which is not part of the training and inference flow.
Table 3: Statistics of datasets.
Dataset # Relation # Entity # Train # Validation # Test
UMLS 46 135 5327 596 633
@@
5 Experiments
5.1 Experiment setting
@@
Data preparation. To evaluate our methods for completing missing triples in KGs, we select three datasets for the KGC task: UMLS [47], WN18RR and FB15k-237 [48]. The statistics of each dataset are presented in Table 3. The UMLS comprises biomedical concepts, including drug and disease names, as well as relations between them, such as diagnosis and treatment. The WN18 and FB15k datasets were initially introduced in the work by [15]. Subsequent research [12, 14], however, revealed that these two datasets are susceptible to test leakage. This issue led to the creation of refined versions: WN18RR and FB15k-237. The textual descriptions of WN18RR and FB15k-237 are sourced from the data made available by KG-BERT [9] and CP-KGC [49], respectively.
@@
During the training process for the task of KGC, each triple $(h,r,t)$ is augmented by creating two queries: $(h,r,?)$ and $(t,r^{\prime},?)$, with the corresponding answers $t$ and $h$. Here, $r^{\prime}$ represents the inverse of $r$. Since obtaining $h_{\text{pred}}$ in Equation 10 necessitates inverse relations, we utilize gpt4-turbo-2024-04-09 to generate these inverses, as outlined in Equation 11. The pairs of $r$ and $r^{\prime}$ for the WN18RR dataset are listed in Table 8.
@@
However, dealing with the FB15k-237 dataset appears to be more complicated. In contrast to the WN18RR dataset, the relations found within FB15k-237 pose a challenge due to their intractability for direct input into a text-based model. To address this concern, we undertake sentence formation of the relations and identify their corresponding inverses with the aid of an LLM. Given the substantial number (237) of relations within the FB15k-237 dataset, we opt to present only a limited subset of these relations in Table 9.
@@
Evaluation Metrics. For each query $(h,r,?)$ or $(t,r^{\prime},?)$, our approach computes a score for each entity, along with determining the rank of the correct answer. The Mean Reciprocal Rank (MRR) and Hit@$k$ values under the filtered protocol [15] are reported based on the computed ranks across all queries. MRR calculates the mean of the reciprocal ranks for the answer entities, while Hit@$k$ determines the percentage of desired entities ranked within the top $k$ positions. The filtered protocol excludes the scores of all known true triples in the training, validation, and testing sets.
@@
Implementation details. To obtain a comprehensive evaluation on the selection of LLMs, we utilize various open-source LLMs to generate predictive descriptions, including Qwen2-7B-Instruct [50], phi-3-small-128k-instruct [51] and phi-3-medium-4k-instruct [51]. In this paper, we refer to these three LLMs as qwen2-7b, phi3-small and phi3-medium for simplicity. A subset of triple examples of WN18RR with textual descriptions and predictive descriptions via phi3-medium is shown in Table 10 (more examples are shown in Table 11 and Table 12). For additional information, please see Appendix A.
Table 4: Knowledge graph completion results for and WN18RR and FB15k-237 datasets. All metrics are in %. The best performance for each metric within each dataset is highlighted in bold, and the best metrics among categories are underlined.
WN18RR FB15k-237 UMLS
MRR Hit@1 Hit@3 Hit@10 MRR Hit@1 Hit@3 Hit@10 MRR Hit@1 Hit@3 Hit@10
NeuralLP [19] 38.1 36.8 38.6 40.8 23.7 17.3 25.9 36.1 48.3 33.2 56.3 77.5
DRUM [20] 38.2 36.9 38.8 41.0 23.8 17.4 26.1 36.4 69.5 54.6 80.8 93.5
RNNLogic [21] 51.3 47.1 53.2 59.7 34.9 25.8 38.5 53.3 74.5 63.0 83.3 92.4
LERP [22] 62.2 59.3 63.4 68.2 - - - - 76.2 64.4 85.5 94.2
TransE [15] 24.3 4.3 44.1 53.2 27.9 19.8 37.6 44.1 - - - 98.9
DistMult [16] 44.4 41.2 47.0 50.4 28.1 19.9 30.1 44.6 39.1 25.6 44.5 66.9
RotatE [17] 47.6 42.8 49.2 57.1 33.8 24.1 37.5 53.3 74.4 63.6 82.2 93.9
ConvE [12] 43.0 40.0 44.0 52.0 32.5 23.7 35.6 50.1 - - - 99.0
TuckER [18] 47.0 44.3 48.2 52.6 35.8 26.6 39.4 54.4 73.2 62.5 81.2 90.0
KG-BERT [9] 21.6 4.1 30.2 52.4 - - - 42.0 - - - 99.0
MTL-KGC [24] 33.1 20.3 38.3 59.7 26.7 17.2 29.8 45.8 - - - -
StAR [10] 40.1 24.3 49.1 70.9 29.6 20.5 32.2 48.2 - - - -
SimKGC [11] 66.6 58.7 71.7 80.0 33.6 24.9 36.2 51.1 79.4 70.4 86.5 94.4
GHN [26] 67.8 59.6 71.9 82.1 33.9 25.1 36.4 51.8 - - - -
CP-KGC [27] 67.3 59.9 72.1 80.4 33.8 25.1 36.5 51.6 - - - -
KICGPT [28] 54.9 47.4 58.5 64.1 41.2 32.7 44.8 55.4 - - - -
KERMIT (qwen2-7b) 69.4 62.0 73.7 83.3 35.1 25.8 38.7 53.7 91.6 84.5 98.6 99.5
KERMIT (phi3-small) 69.5 62.1 73.7 83.0 35.2 25.8 38.8 53.9 92.1 85.6 98.3 99.5
KERMIT (phi3-medium) 70.0 62.9 73.8 83.2 35.9 26.6 39.6 54.7 91.9 85.2 98.6 99.4
5.2 Main results
In this section, we report empirical results of our method, along with those of the chosen algorithms. We adopt the numerical results for NeuralLP, DRUM, RNNLogic, and ConvE as presented in [21]. For LERP, we refer to the values provided in [22]. The metrics for TransE, DistMult, RotatE, TuckER, KG-BERT, MTL-KGC, StAR, and SimKGC are sourced from [11, 27].
As shown in Table 4, our proposed model achieves state-of-the-art performance on the WN18RR and UMLS dataset, surpassing all other methods across all evaluation metrics. Compared to the baseline SimKGC on WN18RR, our model significantly improves MRR (from 66.6% to 70.0%), Hit@1 (from 58.7% to 62.9%), and other key metrics. Similar results can also be observed on UMLS. On the FB15k-237 dataset, KERMIT outperforms other approaches except for KICGPT [28]. The success of KICGPT on FB15k-237 may attribute to the massive demonstrations required by in-context learning used by KICGPT, for athe FB15k-237 dataset exbihits a much denser topological structure.
As presented in Table 4, our proposed model delivers state-of-the-art performance on the WN18RR and UMLS datasets, outperforming all competing methods across all evaluation metrics. On WN18RR, our model significantly surpasses our baseline, SimKGC, achieving an MRR improvement from 66.6% to 70.0% and a Hit@1 increase from 58.7% to 62.9%, alongside other metric enhancements. The UMLS dataset exhibits similar trends, further validating our model’s effectiveness. On the FB15k-237 dataset, KERMIT demonstrates strong performance, outperforming all methods except for KICGPT [28]. The superior results of KICGPT on FB15k-237 may be attributed to its reliance on extensive demonstrations in its in-context learning strategy, particularly suited for the dataset’s much denser topological structure.
@@ 5.2 Main results
Table 5: Knowledge graph completion accuracy of KERMIT and its variants on WN18RR and FB15k-237. "pred-desc": with predictive descriptions; "few-shot": with few-shot triple examples when generating predictive descriptions; "sup": with supervised contrastive loss.
@@
Figure 3: Model performance of text-based methods with the proposed approaches on WN18RR dataset. "w/ inv. rel.": with only inverse relations; "w/ pred. desc": with predictive descriptions. The predictive descriptions are collected using phi3-medium.
5.3 Ablation results
@@
To further demonstrate the effectiveness and generalization ability of predictive descriptions and curated inverse relations, we apply our method to various text-based models, including SimKGC [11], StAR [10], and KG-BERT [9], evaluating their performance on the WN18RR and FB15k-237 datasets (see Figure 3 and 5). The results show a distinct improvement across all metrics after incorporating predictive descriptions and inverse relations, demonstrating our method’s effectiveness as a general enhancement tool for text-based models.
Figure 4: Performance on the WN18RR dataset w.r.t. batch size as achieved by KERMIT. "sup" indicates training with the supervised contrastive strategy, whereas "unsup" represents training with the unsupervised approach.
Table 6: Results given by different tokenization strategies of the query sequence on the WN18RR and FB15k-237 datasets. "sep", "space" and "only $T_{\text{pred}}$" refer to the tokenization strategies specified by Equation 12, 14 and 15, respectively.
@@
Table 7: Predictive descriptions of the triple (cover, derivationally related form, covering) generated by phi3-medium, both with and without few-shot examples in the prompt. Descriptions of the entities involved are available in Table 10.
@@
6 Analysis
6.1 Fine-grained analysis
Supervised contrastive loss. In Figure 4, we present a quantitative analysis of how Hit@1 and Hit@10 are affected by the increasing batch size. When the batch size is relatively small, the scarcity of positive samples causes supervised contrastive training to underperform unsupervised training in terms of Hit@1 metrics. As the batch size grows to 512, however, supervised training reliably exceeds unsupervised performance. This finding underscores the necessity of sufficient positive samples for supervised contrastive training to achieve its optimal accuracy. Furthermore, the markedly superior Hit@10 results for supervised training highlight its robust generalization ability.
@@ 6.1 Fine-grained analysis
$X_{hr}=[x_{\text{CLS}},H_{\text{desc}},x_{\text{SEP}},R_{\text{desc}},x_{\text{SPACE}},T_{\text{pred}},x_{\text{SEP}}]$ (14)
However, since the operation complexity of multi-head attention [52] grows quadratically as the length of the input sequence increases [53], extra information might cause unexpected computational overhead. An alternative version is to represent $X_{hr}$ with just the $T_{\text{pred}}$ to reduce the computational cost, resulting in Equation 15.
$X^{\prime}_{hr}=\left[x_{\text{CLS}},T_{\text{pred}},x_{\text{SEP}}\right]$ (15)
Few-shot triple examples play an important role in the prompting template we use to instruct LLMs, as illustrated in Table 1. We examine the impact of these few-shot examples by comparing the predictive descriptions generated by phi3-medium with and without their inclusion in the prompt, as shown in Table 7. When presented with the query (cover, derivationally related form, ?), incorporating few-shot examples clearly enhances the LLM’s understanding of the task. This results in a more precise and focused predictive description over the unknown answer "covering".
6.2 Case study on selected LLMs
@@ 7 Conclusion
Acknowledgements
This work is supported by the Shandong Key Research and Development Program (No. 2023CXPT065), National Natural Science Foundation of China (No. 62272129) and Taishan Scholars (No. tsqn202408112).
[…]
@@ Appendix A Experiment settings
Comparison of algorithms. In experiment, the performance of our model is compared with that of the following algorithms: (1) Rule learning methods. We have selected a set of algorithms that specialize in learning logical rules for the purpose of completing missing triples. This selection allows for a comprehensive comparison, and the chosen algorithms are NeuralLP [19], DRUM [20], RNNLogic [21] and LERP [22]. (2) Embedding-based methods. Furthermore, we extend our comparison to include various embedding methods, such as TransE [15], DistMult [16], RotatE [17], ConvE [12] and TuckER [18]. These models adopt an approach where representations for entities and relations are learned respectively. (3) Text-based methods. We select the following models that also leverage textual information into our consideration: KG-BERT [9], MTL-KGC [24], StAR [10], SimKGC [11], GHN [26], CP-KGC [27] and KICGPT [28].
Additional implementation details. To ensure a fair comparison with previous studies, we initialize the BERT encoders using pretrained weights from bert-base-uncased (English). The AdamW optimizer [54] with linear learning rate decay is employed to optimize the model parameters. The initial learning rates are set to $8\times 10^{-5}$ for the WN18RR dataset and $1\times 10^{-5}$ for the FB15k-237 dataset. We maintain a fixed batch size of 1024. The temperature parameter $\tau$ is tuned by searching within the set $\{0.02,0.03,0.04\}$, based on results from 65 epochs for WN18RR and 10 epochs for FB15k-237.